OpenStudy (anonymous):

Find the solution of 3 square root x+5=-9, and determine if it is an extraneous solution. The options are: X=4; not extraneous X=4; extraneous X=0; extraneous X=0; not extraneous

OpenStudy (jdoe0001):

so... ok.. what do you get for "x" from \(\bf 3\sqrt{x+5}=-9\)?

OpenStudy (anonymous):

X=4; extraneous ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

is it? what makes you think so?

OpenStudy (jdoe0001):

well, notice \(\bf 3\sqrt{{\color{brown}{ x}}+5}=-9\implies 3\sqrt{{\color{brown}{ 4}}+5}=-9\implies 3\cdot \sqrt{9}=-9 \\ \quad \\ 3\cdot 3\ne -9\) thus, if x = 4, it IS extraneous

OpenStudy (jdoe0001):

even though x = 4 is correct, procedure wise

OpenStudy (anonymous):

I subtracted 5 which got 4. But how do you determine if it is extraneous?